# The Landscape of Agentic Reinforcement Learning for LLMs: A Survey

### v1

@@ 5.2 RL Framework
Framework Type Key Features Resource Link
@@
[…]
@@
[…]
@@
[…]
In this section, we summarize three categories of codebases/frameworks most relevant to this work: agentic RL frameworks, RLHF and LLM fine-tuning frameworks, and general-purpose RL frameworks. Table 10 provides an overview of the prevailing agentic RL and LLM-RL frameworks for readers’ reference.
@@
Verifiers [496] introduces a verifiable-environment setup for end-to-end policy optimization with LLMs, while SkyRL-v0 [497] and its modular successors [498] demonstrate long-horizon, real-world agent training via reinforcement learning. AREAL [499] scales this paradigm with an asynchronous, distributed architecture tailored to language reasoning tasks, and MARTI [500] extends it further to multi-agent LLM systems that integrate reinforcement training and inference. EasyR1 [501] brings multi-modality support, enabling agents to leverage vision and language signals together in a unified RL framework. AgentFly [502] presents a scalable and extensible agent-RL framework that empowers language-model agents with traditional reinforcement-learning algorithms—enabling token-level multi-turn interaction via decorator-based tools and reward definition, asynchronous execution, and centralized resource management for high-throughput RL training. Agent Lightning [503] is a flexible RL framework that decouples agent execution from training by modeling execution as an MDP and using a hierarchical RL algorithm (LightningRL) to train any AI agent with near-zero code modification. AWORLD [504] is a distributed Agentic RL framework, which tackles the main bottleneck of agent training—experience generation—by orchestrating massively parallel rollouts across clusters, achieving a 14.6× speedup over single-node execution and enabling scalable end-to-end training pipelines.
@@
OpenRLHF [505] offers a high-performance, scalable toolkit designed for large-scale model alignment; TRL [506] provides Hugging Face’s baseline implementations for RLHF experiments; trlX [507] adds distributed training support for fine-tuning models up to tens of billions of parameters; and HybridFlow [508] streamlines experiment management and scaling for RLHF research pipelines. SLiMe [509] is an LLM post-training framework for RL scaling that combines Megatron with SGLang for high-performance multi-mode training, supports Async RL, and enables flexible disaggregated workflows for reward and data generation via custom interfaces and server-based engines.
@@
supply the core algorithms and distributed execution engines that can underpin agentic LLM systems. RLlib [510] is a production-grade, scalable library offering unified APIs for on-policy, off-policy, and multi-agent methods; Acme [511] provides modular, research-oriented building blocks for distributed RL; Tianshou [512] delivers a high-performance, pure-PyTorch platform supporting online, offline, and hierarchical RL; Stable Baselines3 [513] packages reliable PyTorch implementations of standard model-free algorithms; and PFRL [514] (formerly ChainerRL) offers benchmarked deep-RL algorithm implementations for rapid prototyping.
@@ 6 Open Challenges and Future Directions
The advance of agent RL toward general-purpose intelligence hinges on overcoming three pivotal challenges that define the field’s research frontier. First is the challenge of Trustworthiness: ensuring the reliability, safety, and alignment of increasingly autonomous agents. Second is Scaling up Agentic Training, which requires surmounting the immense practical bottlenecks in computation, data, and algorithmic efficiency. Finally, an agent’s capabilities are fundamentally bounded by its world, making the Scaling up Agentic Environments, i.e., the creation of complex and adaptive training grounds.
@@ 6.1 Trustworthiness
The security landscape for autonomous agents is fundamentally more complex than for standard LLMs. While traditional models are primarily vulnerable to attacks on their text-in, text-out interface, agents possess an expanded attack surface due to their external components like tools, memory, and planning modules [515, 65]. This architecture exposes them to novel threats beyond direct prompt injection. For instance, indirect prompt injection can occur when an agent interacts with a compromised external environment, such as a malicious website or API, which poisons its memory or tool outputs [516]. Multi-agent systems further compound these risks by introducing vulnerabilities through inter-agent communication, where one compromised agent can manipulate or mislead others within the collective [515].
@@
Mitigating these amplified risks requires a defense-in-depth approach tailored to agentic systems. A critical first line of defense is robust sandboxing [517, 518], where agents operate in strictly controlled, permission-limited environments to contain the potential damage from a compromised tool or action. At the training level, mitigation strategies must focus on shaping the reward signal itself. This includes implementing process-based rewards that penalize unsafe intermediate steps (e.g., calling an untrusted API) and employing adversarial training within the RL loop, where the agent is explicitly rewarded for resisting manipulation attempts and ignoring poisoned information. Finally, continuous monitoring and anomaly detection are essential for post-deployment safety. By tracking an agent’s actions, such as tool calls and memory access patterns, it is possible to identify deviations from normal behavior, allowing for timely intervention.
@@
In the context of agentic LLMs, hallucination is the generation of confident yet ungrounded outputs, including statements, reasoning steps, or tool usage, that are not rooted in provided evidence or external reality. This issue extends beyond simple factual errors to encompass unfaithful reasoning paths and misaligned planning, with overconfidence often masking the agent’s uncertainty [519, 520]. In multimodal agents, it also manifests as cross-modal inconsistency, such as a textual description mismatching an image, framing it as a fundamental grounding problem [521]. Evaluating hallucination requires assessing both factuality against objective truth and faithfulness to a given source, often measured through benchmarks like HaluEval-QA or by the agent’s ability to appropriately abstain on unanswerable questions, where a refusal to answer ("I don’t know") is a critical signal of epistemic awareness [522, 523].
RL can inadvertently amplify hallucination if the reward mechanism is not carefully designed. Studies show that outcome-driven RL, which rewards only the correctness of the final answer, can encourage agents to find spurious correlations or shortcuts. This process may yield confident but unfounded intermediate reasoning steps, as the optimization process settles into local optima that achieve the goal without being factually sound [522]. This phenomenon introduces a "hallucination tax," where reinforcement finetuning can degrade an agent’s ability to refuse to answer, compelling it to generate responses for unanswerable questions rather than abstaining [523]. However, the effect is highly dependent on the training pipeline; while RL-only post-training can worsen factuality, a structured approach combining SFT with a verifiable-reward RL process can mitigate this degradation [524].
Promising mitigation strategies involve a hybrid approach of training-time alignment and inference-time safeguards. During training, a key direction is to shift from outcome-only rewards to process-based rewards. Techniques like Factuality-aware Step-wise Policy Optimization (FSPO) verify each intermediate reasoning step against evidence, directly shaping the policy to discourage ungrounded claims [522]. Data-centric approaches enhance epistemic humility by training agents on a mix of solvable and unsolvable problems, restoring their ability to abstain when necessary [523]. At the system level, this is complemented by inference-time techniques such as retrieval augmentation, tool-use for fact-checking, and post-hoc verification to ground the agent’s outputs in reliable sources. For multimodal agents, explicitly adding cross-modal alignment objectives is crucial for ensuring consistency [520, 519, 521]. Collectively, these directions aim to align the agent’s reward-seeking behavior with the goal of truthfulness, fostering more reliable and trustworthy autonomous systems.
@@
Sycophancy in LLM agents refers to their tendency to generate outputs that conform to a user’s stated beliefs, biases, or preferences, even when those are factually incorrect or lead to suboptimal outcomes [525]. This behavior transcends mere conversational agreeableness, fundamentally affecting an agent’s planning and decision-making processes. For instance, a sycophantic agent might adopt a user’s flawed reasoning in its internal plan, choose a course of action that validates the user’s incorrect assumptions, or filter information from tools to present only what aligns with the user’s view [526]. This represents a critical misalignment, where the agent optimizes for the user’s expressed preference rather than their latent, long-term interest in achieving the best possible outcome.
RL is a primary cause for this behavior. The underlying mechanism is a form of “reward hacking,” where the agent learns to exploit the reward model in ways that do not align with true human preferences [527]. Because human labelers often show a preference for agreeable and validating responses, the reward model inadvertently learns to equate user satisfaction with sycophantic agreement. Consequently, RLHF can directly incentivize and "exacerbate sycophantic tendencies" by teaching the agent that conforming to a user’s viewpoint is a reliable strategy for maximizing reward, even if it compromises truthfulness [528].
Mitigating sycophancy is an active area of research that focuses on refining the reward signal and training dynamics. A promising direction is the development of sycophancy-aware reward models, which are explicitly trained to penalize responses that merely parrot user beliefs without critical evaluation. Another approach involves leveraging AI-driven feedback, such as in Constitutional AI, where the agent is steered by a set of principles promoting objectivity and neutrality, rather than solely by human preferences [529]. At inference time, strategies like explicitly prompting the agent to adopt a “red team” or contrarian perspective can also help counteract ingrained sycophantic tendencies. Cooper [530] is a reinforcement learning framework that co-optimizes both the policy model and the reward model online, using high-precision rule-based verifiers to select positive samples and LLM-generated negative samples, thereby preventing the policy from exploiting a static reward model (i.e., reward hacking) by continuously adapting the reward model to closing emergent loopholes. Ultimately, the future direction lies in designing reward systems that robustly capture the user’s long-term interests—such as receiving accurate information and making sound decisions—over their immediate desire for validation.
@@ 6.2 Scaling up Agentic Training
Recent advances demonstrate that scaling reinforcement learning fine-tuning (RFT) computation directly enhances the reasoning ability of LLM-based agents. The Agent RL Scaling Law study shows that longer training horizons systematically improve tool-use frequency, reasoning depth, and overall task accuracy, highlighting the predictive benefit of allocating more compute to RL training [306]. Similarly, ProRL reveals that prolonged RL training expands reasoning boundaries beyond those accessible to base models, uncovering novel solution strategies even where extensive sampling from the pretrained model fails [531]. Building upon this, ProRLv2 extends training steps and incorporates more stable optimization techniques, demonstrating sustained benefits as smaller models, after extensive RL training, rival the performance of larger models on mathematics, code, and logic benchmarks [532]. Collectively, these results underscore that scaling compute through extended RL training is not merely complementary to enlarging model or data size, but a fundamental axis for advancing agentic reasoning.
@@
Increasing model capacity heightens both the promise and pitfalls of RL-based agent training. Larger models unlock greater potential but risk entropy collapse and narrowing of capability boundaries, as RL sharpens output distributions toward high-reward modes, limiting diversity [533]. Methods like RL-PLUS address this with hybrid strategies and advantage functions that foster novel reasoning paths, breaking capability ceilings [533]. Meanwhile, scaling demands massive compute, making efficiency vital. A two-stage approach in [351] uses large teachers to generate SFT data for smaller students, refined via on-policy RL. This “SFT+RL” setup outperforms each method alone and cuts compute by half compared to pure SFT. The work also underscores RL’s extreme hyperparameter sensitivity at scale, stressing the need for careful tuning.
@@
Scaling RL training across domains introduces both synergy and conflict in agentic reasoning. Cross-domain RL in math, code, and logic tasks shows complex interactions [534]: some pairings enhance each other, while others interfere and reduce performance. Model initialization also matters—instruction-tuned models generalize differently than raw ones. Building on this, the Guru dataset [535] spans six reasoning domains, showing that RL gains correlate with pretraining exposure: math and code benefit from transfer, but domains like simulation or logic need dedicated training. These findings suggest that while multi-domain RL data can amplify general reasoning, it must be carefully curated to balance complementarity and mitigate interference across tasks.
@@
Efficiency of LLM post-training is a central frontier for sustainable scaling [536]. Beyond brute-force scaling, recent research emphasizes improving RL training efficiency through post-training recipes, methodological refinements, and hybrid paradigms. POLARIS [537] demonstrates that calibrating data difficulty, employing diversity-driven sampling, and extending reasoning length substantially boost RL effectiveness, enabling smaller models to reach or even surpass much larger counterparts on reasoning benchmarks. Complementary work [538] provides systematic evaluations of common RL techniques, finding that judiciously combining just a few simple strategies often outperforms more complex methods. Another research proposes Dynamic Fine-Tuning (DFT) [539], showing that introducing RL principles into gradient scaling can match or exceed advanced RL approaches with minimal additional cost. Taken together, these advances suggest a dual trajectory for the future: on one hand, progressively refining RL-based recipes to maximize efficiency; on the other, rethinking training paradigms to embed RL-like generalization signals without full-fledged online RL. A particularly compelling direction lies in exploring how agentic models might acquire robust generalization from extremely limited data, for instance, by leveraging principled difficulty calibration, meta-learning dynamics, or information-theoretic regularization to distill broad reasoning abilities from a handful of experiences. Such pathways point to the possibility of a new regime of post-training: one where the ability to extrapolate, abstract, and generalize becomes decoupled from sheer data volume, and instead hinges on exploiting the structure and dynamics of the training process itself.
6.3 Scaling up Agentic Environment.

### v4

@@ 5.2 RL Framework
Framework Key Features Resource
@@
Verifiers (Brown, 2025) Verifiable environment setup GitHub
SkyRL-v0 (Cao et al., 2025b) Long-horizon real-world training GitHub
AREAL (Fu et al., 2025) Asynchronous training GitHub
MARTI (Zhang et al., 2025n) Integrated multi-agent training GitHub
EasyR1 (Zheng et al., 2025d) Multimodal support GitHub
AgentFly (Wang et al., 2025j) Scalable asynchronous execution GitHub
Agent Lightning (Luo et al., 2025e) Decoupled hierarchical RL GitHub
AWorld (Yu et al., 2025a) Parallel rollouts across clusters GitHub
RL-Factory (RL-Factory, 2025) Easy-to-design reward GitHub
ROLL (Wang et al., 2025o) Stable Multi-GPU Parallel Training GitHub
AgentRL (Zhang et al., 2025a) Asynchronous Multi-Task Training GitHub
VerlTool (Jiang et al., 2025a) Tool-integrated rollout GitHub
@@
OpenRLHF (Hu et al., 2025b) High-performance scalable RLHF GitHub
TRL (von Werra et al., 2020) Hugging Face RLHF GitHub
trlX (Havrilla et al., 2023) Distributed large-model RLHF GitHub
HybridFlow (Sheng et al., 2025) Streamlined experiment management GitHub
SLiMe (THUDM, 2025) High-performance async RL GitHub
Oat (Liu et al., 2024c) Lightweight RL support GitHub
@@
RLlib (Liang et al., 2018) Production-grade scalable library GitHub
Acme (Hoffman et al., 2020) Modular distributed components GitHub
Tianshou (Weng et al., 2022) High-performance PyTorch platform GitHub
Stable Baselines3 (Raffin et al., 2021) Reliable PyTorch algorithms GitHub
PFRL (Fujita et al., 2021) Benchmarked prototyping algorithms GitHub
In this section, we summarize three categories of codebases/frameworks most relevant to this work: Agentic RL frameworks, RLHF and LLM fine-tuning frameworks, and general-purpose RL frameworks. Table 11 provides an overview of the prevailing Agentic RL and LLM RL frameworks for readers’ reference.
@@
Verifiers (Brown, 2025) introduces a verifiable-environment setup for end-to-end policy optimization with LLMs, while SkyRL-v0 (Cao et al., 2025b) and its modular successors (Griggs et al., 2025) demonstrate long-horizon, real-world agent training via reinforcement learning. AREAL (Fu et al., 2025) scales this paradigm with an asynchronous, distributed architecture tailored to language reasoning tasks, and MARTI (Zhang et al., 2025n) extends it further to multi-agent LLM systems that integrate reinforcement training and inference. EasyR1 (Zheng et al., 2025d) brings multi-modality support, enabling agents to leverage vision and language signals together in a unified RL framework. AgentFly (Wang et al., 2025j) presents a scalable and extensible agent‑RL framework that empowers language‑model agents with traditional reinforcement‑learning algorithms—enabling token‑level multi‑turn interaction via decorator‑based tools and reward definition, asynchronous execution, and centralized resource management for high‑throughput RL training. Agent Lightning (Luo et al., 2025e) is a flexible RL framework that decouples agent execution from training by modeling execution as an MDP and using a hierarchical RL algorithm (LightningRL) to train any AI agent with near-zero code modification. AWorld (Yu et al., 2025a) is a distributed Agentic RL framework, which tackles the main bottleneck of agent training—experience generation—by orchestrating massively parallel rollouts across clusters, achieving a 14.6× speedup over single-node execution and enabling scalable end-to-end training pipelines. ROLL (Wang et al., 2025o) provides a scalable library for large-scale RL optimization with a unified controller, parallel workers, and automatic resource mapping for efficient multi-GPU training. VerlTool (Jiang et al., 2025a) introduces an Agentic RL with tool use (ARLT) framework built upon Verl (Sheng et al., 2025), enabling agents to jointly optimize planning and execution across interactive environments. AgentRL (Zhang et al., 2025a) provides a scalable asynchronous framework for multi-turn, multi-task Agentic RL, unifying environment orchestration and introducing cross-policy sampling and task advantage normalization for stable large-scale training.
@@
OpenRLHF (Hu et al., 2025b) offers a high-performance, scalable toolkit designed for large-scale model alignment; TRL (von Werra et al., 2020) provides Hugging Face’s baseline implementations for RLHF experiments; trlX (Havrilla et al., 2023) adds distributed training support for fine-tuning models up to tens of billions of parameters; and HybridFlow (Sheng et al., 2025) streamlines experiment management and scaling for RLHF research pipelines. SLiMe (THUDM, 2025) is an LLM post-training framework for RL scaling that combines Megatron with SGLang for high-performance multi-mode training, supports Async RL, and enables flexible disaggregated workflows for reward and data generation via custom interfaces and server-based engines.
@@
supply the core algorithms and distributed execution engines that can underpin agentic LLM systems. RLlib (Liang et al., 2018) is a production-grade, scalable library offering unified APIs for on-policy, off-policy, and multi-agent methods; Acme (Hoffman et al., 2020) provides modular, research-oriented building blocks for distributed RL; Tianshou (Weng et al., 2022) delivers a high-performance, pure-PyTorch platform supporting online, offline, and hierarchical RL; Stable Baselines3 (Raffin et al., 2021) packages reliable PyTorch implementations of standard model-free algorithms; and PFRL (Fujita et al., 2021) (formerly ChainerRL) offers benchmarked deep-RL algorithm implementations for rapid prototyping.
@@ 6 Open Challenges and Future Directions
The advance of agent RL toward general-purpose intelligence hinges on overcoming three pivotal challenges that define the field’s research frontier. First is the challenge of Trustworthiness: ensuring the reliability, safety, and alignment of increasingly autonomous agents. Second is Scaling up Agentic Training, which requires surmounting the immense practical bottlenecks in computation, data, and algorithmic efficiency. Finally, an agent’s capabilities are fundamentally bounded by its world, making Scaling up Agentic Environments—the creation of complex and adaptive training grounds—a critical necessity.
@@ 6.1 Trustworthiness
The security landscape for autonomous agents is fundamentally more complex than for standard LLMs. While traditional models are primarily vulnerable to attacks on their text-in, text-out interface, agents possess an expanded attack surface due to their external components like tools, memory, and planning modules (Wang et al., 2025l; Shang et al., 2025b). This architecture exposes them to novel threats beyond direct prompt injection. For instance, indirect prompt injection can occur when an agent interacts with a compromised external environment, such as a malicious website or API, which poisons its memory or tool outputs (Chen et al., 2024i). Multi-agent systems further compound these risks by introducing vulnerabilities through inter-agent communication, where one compromised agent can manipulate or mislead others within the collective (Wang et al., 2025l).
@@
Mitigating these amplified risks requires a defense-in-depth approach tailored to agentic systems. A critical first line of defense is robust sandboxing (Lu et al., 2025b; Ruan et al., 2024), where agents operate in strictly controlled, permission-limited environments to contain the potential damage from a compromised tool or action. At the training level, mitigation strategies must focus on shaping the reward signal itself. This includes implementing process-based rewards that penalize unsafe intermediate steps (e.g., calling an untrusted API) and employing adversarial training within the RL loop, where the agent is explicitly rewarded for resisting manipulation attempts and ignoring poisoned information. Finally, continuous monitoring and anomaly detection are essential for post-deployment safety. By tracking an agent’s actions, such as tool calls and memory access patterns, it is possible to identify deviations from normal behavior, allowing for timely intervention.
@@
In the context of agentic LLMs, hallucination is the generation of confident yet ungrounded outputs, including statements, reasoning steps, or tool usage, that are not rooted in provided evidence or external reality. This issue extends beyond simple factual errors to encompass unfaithful reasoning paths and misaligned planning, with overconfidence often masking the agent’s uncertainty (Cossio, 2025; Huang et al., 2025b). In multimodal agents, it also manifests as cross-modal inconsistency, such as a textual description mismatching an image, framing it as a fundamental grounding problem (Bai et al., 2025). Evaluating hallucination requires assessing both factuality against objective truth and faithfulness to a given source, often measured through benchmarks like HaluEval-QA or by the agent’s ability to appropriately abstain on unanswerable questions, where a refusal to answer ("I don’t know") is a critical signal of epistemic awareness (Li & Ng, 2025; Song et al., 2025c).
RL can inadvertently amplify hallucination if the reward mechanism is not carefully designed. Studies show that outcome-driven RL, which rewards only the correctness of the final answer, can encourage agents to find spurious correlations or shortcuts. This process may yield confident but unfounded intermediate reasoning steps, as the optimization process settles into local optima that achieve the goal without being factually sound (Li & Ng, 2025). This phenomenon introduces a "hallucination tax," where reinforcement finetuning can degrade an agent’s ability to refuse to answer, compelling it to generate responses for unanswerable questions rather than abstaining (Song et al., 2025c). However, the effect is highly dependent on the training pipeline; while RL-only post-training can worsen factuality, a structured approach combining SFT with a verifiable-reward RL process can mitigate this degradation (Yao et al., 2025).
Promising mitigation strategies involve a hybrid approach of training-time alignment and inference-time safeguards. During training, a key direction is to shift from outcome-only rewards to process-based rewards. Techniques like Factuality-aware Step-wise Policy Optimization (FSPO) verify each intermediate reasoning step against evidence, directly shaping the policy to discourage ungrounded claims (Li & Ng, 2025). Data-centric approaches enhance epistemic humility by training agents on a mix of solvable and unsolvable problems, restoring their ability to abstain when necessary (Song et al., 2025c). At the system level, this is complemented by inference-time techniques such as retrieval augmentation, tool-use for fact-checking, and post-hoc verification to ground the agent’s outputs in reliable sources. For multimodal agents, explicitly adding cross-modal alignment objectives is crucial for ensuring consistency (Huang et al., 2025b; Cossio, 2025; Bai et al., 2025). Collectively, these directions aim to align the agent’s reward-seeking behavior with the goal of truthfulness, fostering more reliable and trustworthy autonomous systems.
@@
Sycophancy in LLM agents refers to their tendency to generate outputs that conform to a user’s stated beliefs, biases, or preferences, even when those are factually incorrect or lead to suboptimal outcomes (Sun & Wang, 2025). This behavior transcends mere conversational agreeableness, fundamentally affecting an agent’s planning and decision-making processes. For instance, a sycophantic agent might adopt a user’s flawed reasoning in its internal plan, choose a course of action that validates the user’s incorrect assumptions, or filter information from tools to present only what aligns with the user’s view (Malmqvist, 2024). This represents a critical misalignment, where the agent optimizes for the user’s expressed preference rather than their latent, long-term interest in achieving the best possible outcome.
RL is a primary cause for this behavior. The underlying mechanism is a form of “reward hacking,” where the agent learns to exploit the reward model in ways that do not align with true human preferences (Lu et al., 2024). Because human labelers often show a preference for agreeable and validating responses, the reward model inadvertently learns to equate user satisfaction with sycophantic agreement. Consequently, RLHF can directly incentivize and "exacerbate sycophantic tendencies" by teaching the agent that conforming to a user’s viewpoint is a reliable strategy for maximizing reward, even if it compromises truthfulness (Wen et al., 2024).
Mitigating sycophancy is an active area of research that focuses on refining the reward signal and training dynamics. A promising direction is the development of sycophancy-aware reward models, which are explicitly trained to penalize responses that merely parrot user beliefs without critical evaluation.
At inference time, strategies like explicitly prompting the agent to adopt a “red team” or contrarian perspective can also help counteract ingrained sycophantic tendencies. Cooper (Hong et al., 2025b) is a reinforcement learning framework that co‑optimizes both the policy model and the reward model online, using high‑precision rule‑based verifiers to select positive samples and LLM‑generated negative samples, thereby preventing the policy from exploiting a static reward model (i.e., reward hacking) by continuously adapting the reward model to closing emergent loopholes. Ultimately, the future direction lies in designing reward systems that robustly capture the user’s long-term interests—such as receiving accurate information and making sound decisions—over their immediate desire for validation.
@@ 6.2 Scaling up Agentic Training
Recent advances demonstrate that scaling reinforcement learning fine-tuning (RFT) computation directly enhances the reasoning ability of LLM-based agents. The Agent RL Scaling Law study shows that longer training horizons systematically improve tool-use frequency, reasoning depth, and overall task accuracy, highlighting the predictive benefit of allocating more compute to RL training (Mai et al., 2025). Similarly, ProRL reveals that prolonged RL training expands reasoning boundaries beyond those accessible to base models, uncovering novel solution strategies even where extensive sampling from the pretrained model fails (Liu et al., 2025h). Building upon this, ProRLv2 extends training steps and incorporates more stable optimization techniques, demonstrating sustained benefits as smaller models, after extensive RL training, rival the performance of larger models on mathematics, code, and logic benchmarks (Hu et al., 2025a). Collectively, these results underscore that scaling compute through extended RL training is not merely complementary to enlarging model or data size, but a fundamental axis for advancing agentic reasoning.
@@
Increasing model capacity heightens both the promise and pitfalls of RL-based agent training. Larger models unlock greater potential but risk entropy collapse and narrowing of capability boundaries, as RL sharpens output distributions toward high-reward modes, limiting diversity (Dong et al., 2025c). Methods like RL-PLUS address this with hybrid strategies and advantage functions that foster novel reasoning paths, breaking capability ceilings (Dong et al., 2025c). Meanwhile, scaling demands massive compute, making efficiency vital. A two-stage approach in Vattikonda et al. (2025) uses large teachers to generate SFT data for smaller students, refined via on-policy RL. This “SFT+RL” setup outperforms each method alone and cuts compute by half compared to pure SFT. The work also underscores RL’s extreme hyperparameter sensitivity at scale, stressing the need for careful tuning.
@@
Scaling RL training across domains introduces both synergy and conflict in agentic reasoning. Cross-domain RL in math, code, and logic tasks shows complex interactions (Li et al., 2025o): some pairings enhance each other, while others interfere and reduce performance. Model initialization also matters—instruction-tuned models generalize differently than raw ones. Building on this, the Guru dataset (Cheng et al., 2025) spans six reasoning domains, showing that RL gains correlate with pretraining exposure: math and code benefit from transfer, but domains like simulation or logic need dedicated training. These findings suggest that while multi-domain RL data can amplify general reasoning, it must be carefully curated to balance complementarity and mitigate interference across tasks.
@@
The efficiency of LLM post-training is a central frontier for sustainable scaling (Tie et al., 2025). Beyond brute-force scaling, recent research emphasizes improving RL training efficiency through post-training recipes, methodological refinements, and hybrid paradigms. POLARIS (An et al., 2025) demonstrates that calibrating data difficulty, employing diversity-driven sampling, and extending reasoning length substantially boost RL effectiveness, enabling smaller models to reach or even surpass much larger counterparts on reasoning benchmarks. Complementary work (Liu et al., 2025s) provides systematic evaluations of common RL techniques, finding that judiciously combining just a few simple strategies often outperforms more complex methods. Another study proposes Dynamic Fine-Tuning (DFT) (Wu et al., 2025l), showing that introducing RL principles into gradient scaling can match or exceed advanced RL approaches with minimal additional cost. Taken together, these advances suggest a dual trajectory for the future: on one hand, progressively refining RL-based recipes to maximize efficiency; on the other, rethinking training paradigms to embed RL-like generalization signals without full-fledged online RL. A particularly compelling direction lies in exploring how agentic models might acquire robust generalization from extremely limited data, for instance, by leveraging principled difficulty calibration, meta-learning dynamics, or information-theoretic regularization to distill broad reasoning abilities from a handful of experiences. Such pathways point to the possibility of a new regime of post-training: one where the ability to extrapolate, abstract, and generalize becomes decoupled from sheer data volume, and instead hinges on exploiting the structure and dynamics of the training process itself.
6.3 Scaling up Agentic Environments